# Robust LLM Unlearning with MUDMAN: Meta-Unlearning with Disruption Masking And Normalization

### v3

@@ Abstract
Language models can retain dangerous knowledge and skills even after extensive safety fine-tuning, posing both misuse and misalignment risks. Recent studies show that even specialized unlearning methods can be easily reversed. To address this, we systematically evaluate many existing and novel components of unlearning methods and identify ones crucial for irreversible unlearning.
@@
Additionally, we identify the need for normalizing the unlearning gradients, and also confirm the usefulness of meta-learning. We combine these insights into MUDMAN (Meta-Unlearning with Disruption Masking and Normalization) and validate its effectiveness at preventing the recovery of dangerous capabilities. MUDMAN outperforms the prior TAR method by 40%, setting a new state-of-the-art for robust unlearning.
Robust LLM Unlearning with MUDMAN:
Meta-Unlearning with Disruption Masking And Normalization
Filip Sondejthanks: Correspondence to: filipsondej@proton.me Jagiellonian University                        Yushi Yang University of Oxford                        Mikołaj Kniejski University of Warsaw                        Marcel Windys Independent
@@ 1 Introduction
Studies show that popular safety fine-tuning techniques like DPO and RLHF fail to remove dangerous knowledge; instead, they minimally modify model weights to hide it Lee et al. (2024), allowing it to reemerge through jailbreak inputs (Zou et al., 2023), or even accidentally (Qi et al., 2023; Deeb and Roger, 2024). Even specialized unlearning techniques turn out to be easily reversible (Lynch et al., 2024; Łucki et al., 2025; Deeb and Roger, 2024).
To address these challenges, we systematically investigated which components of unlearning algorithms make behavior removal truly irreversible, testing both existing and newly designed ones. We identified several key components for more robust unlearning: Disruption Masking, meta-learning, and gradient normalization. Additionally, we constrain unlearning to specific model modules. We integrate these into MUDMAN (Meta-Unlearning with Disruption Masking and Normalization) and demonstrate that it significantly outperforms state-of-the-art across multiple models and tasks.
Figure 1: MUDMAN pipeline. While training the base model, we periodically fork out an adversary model and train it to perform well on the forget set (orange gradients). Using this adversarial model we calculate the unlearning gradients (red) to be applied to the base model. Before applying the unlearning gradients, we normalize them and perform Disruption Masking, which zeroes out each weight’s unlearning gradient if its sign differs from its retaining gradient (green).
@@ 2 Related Work
To address this, modern unlearning techniques increasingly incorporate MAML (model-agnostic meta-learning) (Tamirisa et al., 2024a; Henderson et al., 2023; Tamirisa et al., 2024b). This approach anticipates how an attacker could relearn the target capability, by deriving unlearning gradients from a copy of the model trained on the forget set Finn et al. (2017).
However, for each existing unlearning method, there are ways to elicit the supposedly removed capabilities, for example with jailbreaks, few-shot prompting, fine-tuning, in-context learning, out-of-distribution inputs or disabling refusal mechanisms with representation engineering (Lynch et al., 2024; Łucki et al., 2025).
3 Methodology
@@
We conduct experiments with three language models of different sizes: pythia-14m (Biderman et al., 2023) [^1] Using pythia-14m allowed us to iterate fast on our techniques. Then we validate our findings on the larger, more modern models., SmolLM-135M (Allal et al., 2024) and Llama-3.2-1B (Grattafiori et al., 2024).
Datasets
@@
Then we move to knowledge unlearning with a realistic target: unlearning biohazardous knowledge using Pile-Bio (Tamirisa et al., 2024a), a subset of the Pile (Gao et al., 2020) containing texts about molecular biology. The rest of the Pile serves as the retain set.
To test unlearning effectiveness on Pile-Bio, we measure the accuracy on WMDP-Bio (Li et al., 2024), a dataset of 1273 multiple-choice questions designed as a proxy measurement of hazardous biosecurity knowledge. We use temperature=0 to get the answers.
@@
Method performance is highly dependent on hyperparameter selection. To ensure a fair comparison, we perform an automated hyperparameter search for each method using Optuna (Akiba et al., 2019). We verify that each search converges and that the chosen hyperparameter ranges are not saturated. Each search includes hundreds of trials, with each trial consisting of an unlearning stage followed by a fixed relearning stage, and the amount of compute in each stage is constant. (See Appendix F for details.)
@@
Optuna tries to maximize forget set loss (or WMDP accuracy) after relearning with supervised fine-tuning. While we are interested in the overall recoverability of unlearned behavior (e.g., through jailbreaks or out-of-distribution attacks), here we just use supervised fine-tuning, as it is the simplest and most reliable way to resurface removed behavior (Lynch et al., 2024).
In addition to maximizing forget set loss, we ensure that unlearning does not significantly degrade performance on the retain set. To control for this, we terminate and reject trials where the retain loss exceeds a fixed threshold [^2] Set as the initial retain loss + 0.05. We add 0.05 to accommodate for random loss fluctuations..
For the WMDP experiments, rather than controlling for the retain loss increase, we use the accuracy on the MMLU benchmark (Hendrycks et al., 2021) to gain a more comprehensive understanding of performance preservation [^3] We exclude all the questions about biology and health. We use temperature=1 to get MMLU answers, to avoid discontinous answer switching which happens with temperature=0, which makes trial rejection less predictable. During unlearning we allow at most 1 percentage point drop in MMLU accuracy..
Figure 2: Ablation study of MUDMAN. To establish that each part of MUDMAN is indeed necessary, we disable them one by one and measure unlearning performance. We also compare to the state-of-the-art TAR method. The baseline is the loss level with no unlearning applied, but after the same relearning as the other methods underwent. Each bar corresponds to one Optuna hyperparameter search. The reported loss is the average of the last 30 valid trials and error bars are their standard error. We can see that Disruption Masking makes a huge difference (orange vs red), and that it accounts for most of the improvement over TAR. Meta-learning and unlearning gradient normalization also tend to help but not in every setup. Sometimes ablating them yields better performance, but insignificantly. In each experiment we only trained the first layers of each MLP.
Figure 3: Accuracy on WMDP-Bio after unlearning and relearning on Pile-Bio. The base level on the rigth (43.5%) is the accuracy with no unlearning applied, but also after relearning. Reported accuracy is the average of the last 60 valid trials in each search. Just like in Figure 2, using Disruption Masking is crucial (orange vs red) and it accounts for most of the improvement over TAR. We also cleary see the need for meta-learning and unlearning gradient normalization. In each experiment we only trained the first layers of each MLP.
@@
1:  ${\color[rgb]{0,0.392,0}retain\_acc}=0$ Initialize retain accumulator
2:  for $loop\_num=1$ to $num\_iterations$ do
@@
4 Building a Robust Unlearning Method
We conducted hundreds of small-scale experiments, testing methods ranging from direct model edits to blocking the updates during relearning. See Appendix D for details of these approaches. While almost all methods succeeded in making the forget set loss high during unlearning, relearning typically restored it immediately.
However, we identified several key components that consistently improve unlearning robustness. We integrate these into MUDMAN (Meta-Unlearning with Disruption Masking and Normalization), which outperforms the state-of-the-art TAR method Tamirisa et al. (2024a). See Algorithm 1 for pseudocode and Appendix E for a minimal PyTorch implementation. Below, we describe each component in detail before presenting MUDMAN’s overall results.
4.1 Meta-Learning
First, we confirm the effectiveness of MAML (model-agnostic meta-learning) (Finn et al., 2017), which is increasingly used in modern unlearning methods (Tamirisa et al., 2024a; Henderson et al., 2023; Tamirisa et al., 2024b). MAML involves training a copy of the main model—which we call the adversary—on the forget set and applying its gradients to the main model.
@@ 4.1 Meta-Learning
Traditional MAML employs an inner loop to train multiple adversaries and accumulate unlearning gradients before updating the main model (Finn et al., 2017). However, in early experiments, we found that interleaving adversary and main model updates improves performance. To achieve this, we flatten the process into a single loop (see Algorithm 1) and focus on training a single adversary deeply, rather than multiple ones briefly.
4.2 Disruption Masking
@@
This is the key contribution of MUDMAN. In existing unlearning techniques (Tamirisa et al., 2024a; Rosati et al., 2024; Li et al., 2024), during unlearning the model is also trained on the retain set, hoping to revert any unintended disruptions. Instead, we take a more selective approach, aiming to avoid disruptions altogether rather than correcting them later.
@@
Our intuition is that since the model has already undergone extensive pre-training, its weights are near-optimal. Any unnecessary modifications risk disrupting well-learned representations, requiring significant compute to recover what pre-training had already established.
@@
To prevent disruption, we found it is best to apply unlearning gradients only when they have the same sign as their respective retain gradients (see Equation 1 and Figure 4). This not only prevents disruption but also passively improves retain performance.
$\begin{matrix}g_{final}=\begin{cases}g_{u}&\text{if }\text{sign}(g_{u})=\text{sign}(g_{r})\\ 0&\text{otherwise}\end{cases}\\ \text{where $g_{u}$ is the unlearning gradient}\\ \text{and $g_{r}$ is the retaining gradient}\end{matrix}$ (1)
@@
To generalize beyond the current batch and identify what might break retain set performance overall, we use SGD momentum (Rumelhart et al., 1986), which adds retain gradients to a decaying accumulator. Then, instead of the signs of the single-batch retain gradients, we use the signs of this accumulator. This incurs no additional memory cost, as we already need to store retain gradients alongside unlearning gradients, so we simply store the accumulator instead.
Normal Disruption Masking
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 4: Disruption Masking and its training dynamics. On the left we show regular unlearning run where we maximize the forget loss while trying to minimize the retain loss. For each weight, we color its unlearning update green if it improves the retain loss, and red if it harms it. Disruption Masking (on the right) filters out all these harmful updates. This way we can raise the forget loss without impacting the retain loss.
4.3 Gradient Normalization
@@
Late in the unlearning process, we observe that gradient norms shrink significantly, causing unlearning to slow down. To address this, we normalize the unlearning gradients so that the strength of unlearning stays constant.
Concretely, we divide each unlearning gradient by a global gradient norm calculated as: $(\sum_{m\in\text{modules}}\|unlearning\_grad_{m}\|^{2})^{0.5}$.
@@
This not only improves overall performance but also makes the process easier to tune. See Appendix B for a comparison of various normalization variants.
4.4 Module Selection
@@
We found that it helps to also be selective about which modules of the model we intervene on. For example MLP’s second weight matrices and attention’s Q and K matrices tend to disrupt general performance a lot, while not being that important for unlearning. The most effective unlearning targets turn out to be the first layers of MLPs (and in the case of gated MLPs, the gating matrices). This is likely because only these modules are able to deactivate MLP neurons, and once a neuron is inactive, neither its incoming nor outgoing weights can be updated by backpropagation, effectively preventing relearning. See Appendix A for detailed comparisons of unlearning on various modules.
4.5 Experimental Results
@@
We validated MUDMAN on three models and two unlearning tasks. For a clean comparison with TAR (Tamirisa et al., 2024a), we adapted its implementation to match our setup: using a single training loop and omitting its initial representation noising step. As shown in Figure 2, MUDMAN consistently outperforms this adapted TAR across all tested cases. In individual setups some ablations occasionally match full MUDMAN, but never significantly surpass it.
MUDMAN is compatible with any unlearning loss. In Figure 2, we used negative cross-entropy loss as it performed best. However, in the WMDP setup shown on Figure 3, MUDMAN achieves the strongest results with negative entropy loss (Tamirisa et al., 2024a), so we adopted it there.
@@ 5 Conclusion
We introduce MUDMAN, a novel method consisting of three components: meta-unlearning, Disruption Masking and normalization of unlearning gradients. We have shown across multiple datasets and models that each of these components significantly improves unlearning robustness. They also come at minimal computational and memory overhead.
Our results demonstrate the success of selective unlearning methods which do not disrupt model performance, inviting future work to pursue this line of research.
These advances bring us closer to the goal of truly irreversible unlearning, which is critical for safe deployment as language models continue to acquire dangerous knowledge and capabilities.
Acknowledgments
We thank Stephen Casper, Kay Kozaronek and Artyom Karpov for valuable discussions and feedback.
@@ 6 Limitations
Elicitation Methods
A limitation of current study is that we only used supervised fine-tuning to elicit unwanted behavior. While it is the most reliable elicitation method (Lynch et al., 2024), it could be complemented by others, such as jailbreaks, few-shot prompting, in-context learning, out-of-distribution attacks, and automated curiosity-based red teaming (Zheng et al., 2025; Hong et al., 2024) to test unlearning robustness.
Additionally, to settle whether our techniques truly remove capabilities or just make them harder to recover, one could use the approach introduced by Deeb and Roger (2024), where the attacker tries to uncover some unknown, supposedly unlearned facts, by fine-tuning on another, non-overlapping set of facts.
@@
For simplicity we only evaluated monolithic methods which apply a single algorithm throughout the whole unlearning process. This involved stripping our TAR baseline (Tamirisa et al., 2024a) of its initial representation noising step and only keeping the meta-learning core of the algorithm. Future work could investigate the effectiveness of consecutively applying different unlearning methods, looking for synergies between them.
Selectivity and Granularity
In our experiments, we have generally seen good results from techniques which aim to be more selective and granular. For instance, we think that trying to understand what happens at the level of individual logits and tweaking unlearning loss functions to target only the crucial logits is a ripe area for future improvements. Additionally, a more granular, per-token analysis of forget set loss could be valuable - rather than concentrating the loss increase on a few unwanted tokens, the goal should be a high loss across all unwanted tokens. Further work should also try to pin down with more certainty which modules in a model are optimal to target for unlearning.
Appendix A Target Modules

### v4

# Biosecurity Safeguards for Generative AI

@@ Abstract
Language models can retain dangerous knowledge and skills even after extensive safety fine-tuning, posing both misuse and misalignment risks. Recent studies show that even specialized unlearning methods can be easily reversed. To address this, we systematically evaluate existing ones and propose novel components of unlearning methods and identify ones crucial for irreversible unlearning.
@@
Additionally, we identify the need for normalizing the unlearning gradients, and also confirm the usefulness of meta-learning. We combine these insights into MUDMAN (Meta-Unlearning with Disruption Masking and Normalization) and validate its effectiveness at preventing the recovery of dangerous capabilities. MUDMAN outperforms the prior TAR method by 40%, setting a new state-of-the-art for robust unlearning. Code: anonymous.4open.science/r/MUDMAN
@@ 1 Introduction
Studies show that popular safety fine-tuning techniques like DPO and RLHF fail to remove dangerous knowledge; instead, they minimally modify model weights to hide it (Lee et al., 2024), allowing it to reemerge through jailbreak inputs (Zou et al., 2023), or even accidentally (Qi et al., 2023; Deeb and Roger, 2024). Even specialized unlearning techniques turn out to be easily reversible (Lynch et al., 2024; Łucki et al., 2025; Deeb and Roger, 2024). It remains unclear which components of unlearning algorithms can make harmful capabilities truly irreversible.
To address these challenges, we systematically investigate which components of unlearning algorithms make behavior removal truly irreversible, testing both existing and proposing newly designed ones. We identify several key components for more robust unlearning: Disruption Masking, meta-learning, and gradient normalization. We integrate these into MUDMAN (Meta-Unlearning with Disruption Masking and Normalization) and demonstrate that it significantly outperforms state-of-the-art across multiple models and tasks.
Figure 1: The MUDMAN pipeline. When training the base model, we periodically fork out an adversary model and train it to perform well on the forget set (orange gradients). Using this adversarial model, we calculate the unlearning gradients (red) to be applied to the base model. Before applying the unlearning gradients, we normalize them and perform Disruption Masking, which zeroes out each weight’s unlearning gradient if its sign differs from its retaining gradient (green).
Normal Disruption Masking
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Disruption Masking and its training dynamics. Left: regular unlearning runs where we maximize the forget loss while trying to minimize the retain loss. For each weight, we color its unlearning update green if it improves the retain loss, and red if it harms it. Right: Disruption Masking filters out all these harmful updates. This manages to raise the forget loss without impacting the retain loss.
@@ 2 Related work
To address this, modern unlearning techniques increasingly incorporate MAML (model-agnostic meta-learning) (Tamirisa et al., 2024a; Henderson et al., 2023; Tamirisa et al., 2024b). This approach anticipates how an attacker could relearn the target capability, by deriving unlearning gradients from a copy of the model trained on the forget set (Finn et al., 2017).
However, for each existing unlearning methods, there are still ways to elicit the supposedly removed capabilities, such as jailbreak prompts, few-shot prompting, fine-tuning, in-context learning, out-of-distribution inputs or disabling refusal mechanisms with representation engineering (Lynch et al., 2024; Łucki et al., 2025).
3 Experiment setup
@@
We conduct experiments with three language models of different sizes: pythia-14m (Biderman et al., 2023). SmolLM-135M (Allal et al., 2024) and Llama-3.2-1B (Grattafiori et al., 2024). Using pythia-14m allowed us to iterate fast on our techniques, then we validate our findings on the larger, more modern models.
Unlearning datasets
@@
Then we move to knowledge unlearning with a realistic harmful target: unlearning biohazardous knowledge using Pile-Bio (Tamirisa et al., 2024a), a subset of the Pile (Gao et al., 2020) containing texts about molecular biology. The rest of the Pile serves as the retain set.
To test unlearning effectiveness on Pile-Bio, we measure the accuracy on WMDP-Bio (Li et al., 2024), a dataset of 1273 multiple-choice questions designed as a proxy measurement of hazardous biosecurity knowledge. We use temperature=0 to generate the answers.
@@
Method performance is highly dependent on hyperparameter selection. To ensure a fair comparison, we perform an automated hyperparameter search for each tested method using Optuna (Akiba et al., 2019). We verify that each search converges and that the chosen hyperparameter ranges are not saturated. Each search includes hundreds of trials, with each trial consisting of an unlearning stage followed by a fixed relearning stage, and the amount of compute in each stage is constant. (See Appendix G for details.)
@@
Optuna tries to maximize forget set loss (or WMDP accuracy) after relearning with supervised fine-tuning. While we are interested in the overall recoverability of unlearned behavior (e.g., through jailbreaks or out-of-distribution attacks), here we focus on supervised fine-tuning, as it is the most straightforward and most reliable way to resurface removed behavior (Lynch et al., 2024).
In addition to maximizing forget set loss, we ensure that unlearning does not significantly degrade performance on the retain set. To control for this, we terminate and reject trials where the retain loss exceeds a fixed threshold. [^1] Set as the initial retain loss + 0.05. We add 0.05 to accommodate for random loss fluctuations.
For the WMDP experiments, rather than controlling for the retain loss increase, we use the accuracy on the MMLU benchmark (Hendrycks et al., 2021) to gain a comprehensive understanding of performance preservation.  [^2] We use temperature=1 to get MMLU answers, to avoid discontinuous answer switching which happens with temperature=0, which makes trial rejection less predictable. During unlearning we allow at most 1 percentage point drop in MMLU accuracy.
Figure 3: Ablation study of MUDMAN. To establish that each part of MUDMAN is indeed necessary, we disable them one by one and measure unlearning performance in terms of the forget loss. We also compare to the state-of-the-art TAR method. The baseline is the loss level with no unlearning applied, but after the same relearning as the other methods underwent. Each bar corresponds to one Optuna hyperparameter search. The reported loss is the average of the last 30 valid trials in that search and error bars are their standard error. It shows that Disruption Masking makes a huge difference (orange vs red), and that it accounts for most of the improvement over TAR. Meta-learning and unlearning gradient normalization also help but not in every setup. Sometimes ablating them yields better performance, but insignificantly. In MUDMAN we use negative cross-entropy as the unlearning loss, due to its top performance. As discussed in Section 4.4, in each experiment we focus on training the first layers of each MLP.
Figure 4: Accuracy on WMDP-Bio after unlearning and relearning on Pile-Bio. The base level on the right (45.5%) is the accuracy with no unlearning applied, but after relearning. Reported accuracy is the average of the last 60 valid trials in each hyperparameter search. We only show Llama, because smaller models had near-random accuracy on WMDP-Bio. Here, we use MUDMAN with negative entropy (Tamirisa et al., 2024a) as the unlearning loss. Similarly to Figure 3, using Disruption Masking is crucial (orange vs red) and it accounts for most of the improvement over TAR. There is also clear improvements from meta-learning and unlearning gradient normalization.
@@
[…]
2:for loop_num=1loop\_num=1 to num_iterationsnum\_iterations do
@@
4 Building a robust unlearning pipeline
We conducted hundreds of small-scale experiments, testing methods ranging from direct model edits to blocking the updates during relearning. See Appendix E for details of approaches. While almost all methods succeeded in making the forget set loss high during unlearning, relearning typically restored it immediately.
Among them, we identify several key components that consistently improve unlearning robustness. We also propose a novel method called Disruption Masking. We integrate these into MUDMAN (Meta-Unlearning with Disruption Masking and Normalization), which outperforms the state-of-the-art TAR method Tamirisa et al. (2024a). See Algorithm 1 for pseudocode and Appendix F for a PyTorch implementation. Below, we describe each component of MUDMAN and its overall results.
4.1 Meta-learning is effective
First, we confirm the effectiveness of MAML (model-agnostic meta-learning) (Finn et al., 2017), which is increasingly used in modern unlearning methods (Tamirisa et al., 2024a; Henderson et al., 2023; Tamirisa et al., 2024b). MAML involves training a copy of the main model, which we call the adversary, on the forget set and applying its gradients to the main model.
@@ 4.1 Meta-learning is effective
Traditional MAML employs an inner loop to train multiple adversaries and accumulate unlearning gradients before updating the main model (Finn et al., 2017). However, in early experiments, we found that interleaving adversary and main model updates actually improves performance. Therefore, we flatten the process into a single loop (see Algorithm 1) and focus on training a single adversary deeply, rather than multiple ones briefly.
4.2 Disruption Masking drives performance
@@
This is our key contribution from MUDMAN. In existing unlearning techniques (Tamirisa et al., 2024a; Rosati et al., 2024; Li et al., 2024), the model is also trained on the retain set during unlearning, hoping to revert any unintended disruptions. Instead, we adopt a more selective approach, aiming to avoid disruptions altogether rather than correcting them later.
@@
The intuition for such high selectivity is that since the model has already undergone extensive pre-training, its weights are near-optimal. Any unnecessary modifications risk disrupting well-learned representations, requiring significant compute to recover what pre-training had already established.
@@
We formalise disruption masking as follows: we apply unlearning gradients only when they have the same sign as their respective retain gradients (see Equation 1 and Figure 2). This not only prevents disruption but also passively improves the retain performance.
[…]
@@
Implementation-wise, to generalize beyond the current batch and identify what might break retain set performance overall, we use SGD momentum (Rumelhart et al., 1986), which adds retain gradients to a decaying accumulator. Then, instead of the signs of the single-batch retain gradients, we use the signs of this accumulator. This incurs no additional memory cost, as we already need to store retain gradients alongside unlearning gradients, so we simply store the accumulator instead.
4.3 Gradient normalization improves speed and stability
@@
In the later phase of the unlearning process, we observe that gradient norms shrink significantly, causing unlearning to slow down. To address this, we normalize the unlearning gradients so that the strength of unlearning remains constant. Specifically, we divide each unlearning gradient by a global gradient norm, calculated as: $(\sum_{m\in\text{modules}}\|unlearning\_grad_{m}\|^{2})^{0.5}$.
@@
This not only improves overall performance but also makes the training process easier to tune. See Appendix C for a comparison of other normalization variants.
4.4 Module selection improves selectivity
@@
We selectively filter out model modules and test which ones to intervene on for more selective unlearning. On Figure 5 we see that MLP’s second weight matrices and attention’s Q and K matrices tend to significantly disrupt general performance, while not being very important for unlearning performance. The most effective unlearning targets are the first layers of MLPs, and in the case of gated MLPs, the gating matrices. This is likely because only these modules are able to deactivate MLP neurons, and once a neuron is inactive, neither its incoming nor outgoing weights can be updated by backpropagation, effectively preventing relearning. See Appendix B for detailed comparisons of unlearning on various modules.
4.5 Improvements from MUDMAN
@@
We validate the superiority of MUDMAN on three models and two unlearning tasks over TAR (Tamirisa et al., 2024a), a state-of-the-art unlearning method. For a clean comparison, we adapt the TAR implementation to match our setup: using a single training loop and omitting its initial representation noising step. As shown in Figure 3, MUDMAN consistently outperforms this adapted TAR across all tested cases. In individual setups, while some ablations occasionally match the full MUDMAN, none of them ever significantly surpasses it.
@@ 5 Conclusion
We introduce MUDMAN, a novel unlearning pipeline consisting of three components: meta-unlearning, Disruption Masking and normalization of unlearning gradients. We show across multiple datasets and models that each of these components significantly improves the robustness of unlearning of harmful knowledge. They also come at minimal computational and memory overhead and are compatible with any unlearning loss.
Our results demonstrate the success of selective unlearning methods which do not disrupt model performance in the first place, inspiring future work to pursue this line of research. Our findings bring closer to the goal of truly irreversible unlearning, which is critical for safe deployment as language models as they continue to acquire dangerous knowledge and capabilities.
@@ 6 Limitations
More elicitation methods
In this study, we only use supervised fine-tuning to elicit unwanted behavior. While it is the most powerful and reliable elicitation method (Lynch et al., 2024), future work could extend to evaluate other attacks, in which the attacker does not have weight access, for example with jailbreak prompts (Zou et al., 2023).
To further test whether our techniques truly remove capabilities or just make them harder to recover, one could use the approach introduced by Deeb and Roger (2024), where the attacker tries to uncover some unknown, supposedly unlearned facts, by fine-tuning on another, non-overlapping set of facts.
@@
For simplicity, we only evaluate monolithic methods which apply a single algorithm throughout the whole unlearning process. This involved stripping our TAR baseline (Tamirisa et al., 2024a) of its initial representation noising step and only keeping the meta-learning core of the algorithm. Future work could investigate the effectiveness of consecutively stacking different unlearning methods, looking for synergies between them.
More work on selectivity and granularity
In our experiments, we generally see good results from techniques aiming to be more selective and granular. Future work could try to understand what happens at the level of individual logits and tweaking unlearning loss functions to target only the crucial logits. Additionally, a more granular, per-token analysis of forget set loss could be valuable: rather than concentrating the loss increase on a few unwanted tokens, the goal could be set to a high loss across all unwanted tokens.
Acknowledgments and Disclosure of Funding
We thank Fabien Roger, Stephen Casper, Adam Mahdi, Kay Kozaronek and Artyom Karpov for valuable discussions and feedback.
Filip Sondej’s work was partially funded by a grant from Open Philanthropy.
Appendix A Future work
Dataset design
The spirit of making unlearning more selective can be also applied to dataset design. Usually, the bulk of the forget set is made of benign tokens (filler words, irrelevant passages, etc.), and we are only interested in breaking a smaller subset of critical tokens. So we should find a scalable way to label the importance of each token in the forget sets we use. This would aid both in unlearning that does not disrupt the model, and in a more accurate evaluation of unlearning methods.
Interpretability
We believe that a lot of insight can be extracted from using interpretability on the effects of unlearning techniques. One thing we find particularly mysterious, is that during relearning we often find that forget loss decrease is stopped at some high level and can stay like this for a long time. Then it sometimes suddenly "breaks through", which is reminiscent of grokking. We would like to understand what exactly happens around this "break through", and what mechanistically differentiates cases where this happens from ones where it does not. Also, if we disallow fine-tuning, are the models before "break through" more resistant to other elicitation techniques? Interpretability could start by identifying which neurons are affected by unlearning the most, how their triggering context is shifted and how they relate to other neurons.
Adversary depth
Unlearning can happen at different "depths", defined as the length of training of the adversary. Unlearning with shallow adversaries or no adversary at all tends to be easily reverted, while more depth can be too computationally costly. It would be illuminating to check in isolation what happens at various depths.
Relation to deliberative alignment
Finally, the concept of unlearning could be inverted, and we could try to make sure that some critical desired behavior always happens, and will never be disturbed out-of-distribution or under jailbreaks or fine-tuning. This is particularly useful in the deliberative alignment framework [Guan et al., 2025], where we want some lines of reasoning to trigger reliably in certain contexts.
Appendix B Target modules
Figure 5: Performance comparison across different target module configurations for unlearning. Higher values indicate better unlearning effectiveness while maintaining model capabilities. The baseline is the loss without any unlearning, but with the same relearning stage as all the methods underwent. Gate projection (and in the case of pythia its equivalent–the first MLP layer) helps most consistently. Other potential candidates for intervention are V, O and up projections. Q, K and down projections disrupt retain performance so much that is it better to omit them. In case of Pythia, Q, K and V matrices are integrated into one module, so we were not able to analyze them in separation.